OpenStudy (anonymous):

Subtract: x^2/x-6 - 6x/x-6

OpenStudy (anonymous):

\[\frac{ x^{2} }{x-6? } - \frac{ 6x }{ x-6}\]

OpenStudy (anonymous):

The denominator is common. hence - x^2-6x/x-6 => x(x-6)/x-6.

OpenStudy (anonymous):

So I multiply both numerators by the denominator?

OpenStudy (anonymous):

@luckythebest

OpenStudy (anonymous):

Nope. you don't do that. The thing is, since the denominator is the same, \[\frac{ x^2 }{ x-6 } - \frac{ 6x }{ x-6 }\]

OpenStudy (anonymous):

instead of doing this - you can do x^2 - 6x \[\frac{ x^2-6x }{ x-6 }\]

OpenStudy (anonymous):

\[\frac{ x(x-6) }{ x-6 }\]

OpenStudy (anonymous):

So do I cross out?

OpenStudy (anonymous):

@Ozzy733 yep. the answer is X

OpenStudy (anonymous):

Great, thanks so much!